OpenStudy (anonymous):

using exact value show that sinx/cosx=tanx , x=30 degrees

OpenStudy (anonymous):

sin(30) = 0.5 and cos(30) = sqrt3/2 thus 0.5/sqrt3/2 = sqrt3/3 tan(30) = sqrt3/3

OpenStudy (anonymous):

sinx/cosx is tanx. so any value would be equal because they are the same. Also sinx is opp/hyp and cosx is adj/hyp so if you divide the two you end up with opp/adj which is tanx. Idk if you already knew that or not but that is why they are equal and how I think of it if I forget. sin(pi/6)=1/2 cos(pi/6)=sqrt3/2 so sin(pi/6)/cos(pi/6)= (1/2)/(sqrt3/2)=1/sqrt3

OpenStudy (anonymous):

answer should be square 3 / 3 though

OpenStudy (anonymous):

that is what i gave you.

OpenStudy (anonymous):

oh okay lol thanks

OpenStudy (anonymous):

sqrt= square root

OpenStudy (anonymous):

The answer I gave you is before it is rationalized. Some teachers care if it is some dont.